# Automatic Pruning via Structured Lasso with Class-wise Information

### v1

@@ Abstract
Most pruning methods concentrate on unimportant filters of neural networks. However, they face the loss of statistical information due to a lack of consideration for class-wise data. In this paper, from the perspective of leveraging precise class-wise information for model pruning, we utilize structured lasso with guidance from Information Bottleneck theory. Our approach ensures that statistical information is retained during the pruning process. With these techniques, we introduce two innovative adaptive network pruning schemes: sparse graph-structured lasso pruning with Information Bottleneck (sGLP-IB) and sparse tree-guided lasso pruning with Information Bottleneck (sTLP-IB). The key aspect is pruning model filters using sGLP-IB and sTLP-IB to better capture class-wise relatedness. Compared to multiple state-of-the-art methods, our approaches demonstrate superior performance across three datasets and six model architectures in extensive experiments. For instance, using the VGG16 model on the CIFAR-10 dataset, we achieve a parameter reduction of 85%, a decrease in FLOPs by 61%, and maintain an accuracy of 94.10% (0.14% higher than the original model); we reduce the parameters by 55% with the accuracy at 76.12% using the ResNet architecture on ImageNet (only drops 0.03%). In summary, we successfully reduce model size and computational resource usage while maintaining accuracy. Our codes are at https://anonymous.4open.science/r/IJCAI-8104.
1 Introduction
@@
Deep Learning Krizhevsky et al. (2012), especially CNN-based methods Simonyan and Zisserman (2014); Szegedy et al. (2015); He et al. (2016), has proven its effectiveness in tasks like object detection and classification. Recently, researchers have been attempting to deploy neural network models on resource-constrained edge devices for real-time analysis Liu et al. (2024c); Yu et al. (2024). However, these network parameters are often too large, making deployment difficult. For instance, the VGGNet-16 model has parameters exceeding 500MB Simonyan and Zisserman (2014), which introduces a heavy memory burden and computation overhead.
@@
Fortunately, scientists have discovered that the parameters are usually redundant and the key weights of neural networks exhibit sparsity characteristics He et al. (2017). As a result, a significant amount of work has focused on pruning unimportant weights to reduce the size of the model, which reduces the model’s FLOPs and lowers computational overhead as well. The common methods can be primarily divided into two categories: unstructured pruning and structured pruning Shao and Shin (2022). Unstructured pruning mainly focuses on weight pruning Ding et al. (2019) and N:M sparsity Zhang et al. (2022). However, both methods require hardware support, making them less scalable. Therefore, this paper primarily focuses on structured pruning, which benefits from removing unimportant filters within the CNN.
@@
To avoid the limitations of hand-crafted approaches, most structured pruning methods utilize statistical information Zhang et al. (2021); Barisin and Horenko (2024) to analyze filters, with the aim of minimizing the loss before and after pruning and achieving better accuracy results. However, these methods all overlook the class-wise information Hu et al. (2016), resulting in a less precise pruning scheme. Furthermore, an ideal pruning method should also include an adaptive approach Guo et al. (2023); Xie et al. (2024) to control the level of sparsity based on specific requirements.
Figure 1: The overview of our methods. To utilize information bottleneck theory, ① reshape the input and output feature maps. ② transforms the reshaped feature maps with Gram matrix. Then, we utilize the structured lasso to account for the class-wise information to get the linkage matrix and prune the filters.
@@
Based on these, we employ the precise Information Bottleneck (IB) Tishby et al. (2000) theory perspective to model statistical information of feature maps of CNN for pruning, aiming to minimize the loss from pruning. We also utilize structured lasso techniques that account for sparsity, such as graph-structured lasso Liu et al. (2024b) and tree-guided lasso Kim and Xing (2009); Liu et al. (2024a), which can consider the relatedness in feature maps and form corresponding sub-groups and sub-trees. This approach helps to incorporate class-wise information. Consequently, we propose two novel methods: Sparse Graph-structured Lasso Pruning with Information Bottleneck (sGLP-IB) and Sparse Tree-guided Lasso Pruning with Information Bottleneck (sTLP-IB). Our methods perform pruning at the layer level rather than end-to-end, ensuring more precise pruning filters and preventing cumulative errors layer by layer Sakamoto and Sato (2024). Based on this, our achievements have already been surprising as we implement structured lasso into pruning. Further, we (1) combine Gram matrix Lanckriet et al. (2002) to reduce the dimensions and accelerate computation without compromising accuracy, (2) propose an adaptive method to ensure that our method can be automatically suitable for varying sparsity requirements rather than setting hyperparameters based on experiences, (3) use novel optimization method which has polynomial time complexity with guaranteed convergence rate. Our main contributions are summarized as follows:
@@
  • We are the first to discover pruning with structured lasso from the information bottleneck perspective via precise class-wise information. By enhancing techniques such as graph-structured lasso and tree-guided lasso with information bottleneck, we propose two novel methods: sGLP-IB and sTLP-IB.
@@
  • We have optimized our proposed methods, sGLP-IB and sTLP-IB, to ensure faster computation using Gram matrix and adaptively address different model sizes or FLOPs requirements with our proposed algorithm. These optimizations still guarantee convergence and polynomial computation time complexity using the proximal gradient descent method.
@@
  • We conduct comparative experiments against numerous baselines using two real-world datasets, CIFAR-10, CIFAR-100 and ImageNet. We tested multiple variants of VGGNet, ResNet, and GoogleNet. The experimental results demonstrate the effectiveness of our methods across three metrics: accuracy, memory size, and FLOPs. They both achieve state-of-the-art results, while sTLP-IB demonstrates a stronger ability to learn class information compared to sGLP-IB. Extensive ablation studies further confirm the robustness of our methods and the potential of utilizing class-wise information.
2 Related Works
2.0.1 Structured Pruning
@@
MPF He et al. (2019) considers the geometric distance between filters and neighboring filters to guide the pruning process. CPMC Yan et al. (2021) leverages relationships between filters in the current layer and subsequent layers for pruning decisions. HRank Lin et al. (2020) compresses models by constructing low-rank matrices using information from the same batch. Scop Tang et al. (2020) reduces filters based on their deviation from the expected network behavior. GDP Guo et al. (2021) employs gates with differentiable polarization, learning whether gates are zero during training for pruning purposes. EEMC Zhang et al. (2021) uses a multivariate Bernoulli distribution along with stochastic gradient Hamiltonian Monte Carlo for pruning. SRR-GR Wang et al. (2021) identifies redundant structures within CNNs. FTWT Elkerdawy et al. (2022) predicts pruning strategies using a self-supervised mask. DECORE Alwani et al. (2022) assigns an agent to each filter and uses lightweight learning to decide whether to keep or discard each filter. EPruner Lin et al. (2021) employs Affinity Propagation for efficient pruning. DPFPS Ruan et al. (2021) directly learns the network with structural sparsity for pruning. CC Li et al. (2021) combines tensor decomposition for filter pruning. NPPM Gao et al. (2021) trains the network to predict the performance of the pruned model, guiding the pruning process. LRF-60 Joo et al. (2021) uses Linearly Replaceable Filters to aid pruning decisions. AutoBot Castells and Yeom (2021) tests each filter one-by-one to ensure pruning precision. PGMPF Cai et al. (2022) utilizes prior masks as the basis for pruning masks. DLRFC He et al. (2022) uses Receptive Field Criterion to measure filter importance. FSM Duan et al. (2022) aggregates feature and filter information to evaluate their relevance for shifting. Rrandom Li et al. (2022) employs random search strategies for pruning. The study Hussien et al. (2024) employs activation statistics, foundations in information theory, and statistical analysis with designed regularizations for pruning. CSD Xie et al. (2024) uses channel spatial dependability metric and leverages feature characteristics for pruning. However, the statistical learning information considered by these methods may not be precise enough.
2.0.2 Information Bottleneck and Lasso Principle
@@
The Information Bottleneck (IB) Tishby et al. (2000) method extracts the most relevant output information by considering the input. Aside from the previous mentioned methods, the works Tishby and Zaslavsky (2015); Dai et al. (2018) were the first to use IB for compressing CNNs. FPGM He et al. (2018) employs the geometric median for pruning. L1 Li et al. (2016) illustrates the correspondence between Lasso and IB for pruning. The approach in He et al. (2017) utilizes a two-step optimization process incorporating Lasso for pruning. Papers Scardapane et al. (2016, 2016) leverage the $l_{2}$ lasso (e.g., group lasso) for pruning, with analyzed complexity Rajaraman et al. (2023). APIB Guo et al. (2023) integrates IB with the Hilbert-Schmidt Independence Criterion Lasso for exploration. The paper Sakamoto and Sato (2024) also shows through analysis that layer-level pruning, based on the Hilbert-Schmidt independence criterion, is preferable to end-to-end pruning.
@@
However, none of these methods account for class-wise information. To address this, from the information bottleneck perspective, we employ state-of-the-art techniques like graph-structured lasso Liu et al. (2024b) and tree-guided lasso Liu et al. (2024a) to aggregate class-wise information effectively to propose our novel methods and prune filters inside CNNs.
3 Our Method
3.1 Design Overview
@@
The overview of our methods is shown in Figure 1. To avoid the imprecision that end-to-end methods might introduce, we perform layer-wise pruning independently Sakamoto and Sato (2024). We reshape each layer’s input and output feature map tensors into two-dimensional matrices, enabling us to apply the information bottleneck theory to determine which filters can be pruned. To further reduce computational complexity, capture internal geometric structures, and utilize more flexible kernel functions, we employ Gram matrix Lanckriet et al. (2002). Since lasso aligns well with the information bottleneck theory Tishby and Zaslavsky (2015); Dai et al. (2018), we use structured lasso for solving this problem. This approach also takes into account class-wise information in the feature maps (as well as filters), resulting in a sparse linkage map between layers’ feature maps for pruning.
@@
Math Notations: In this paper, given a CNN model $M$ with $N$ layers, we let $C_{i}$ denote the $i$-th layer of $M$. The input feature map of $C_{i}$ is $X_{i}$ with the shape $(bs,h_{i},w_{i},in_{i})$; the output feature map of $C_{i}$ is $X_{i+1}$ with the shape $(bs,h_{i+1},w_{i+1},out_{i})$, where $i\in\{1,2,...,N\}$, $X_{1}$ is the input sample batch from the datasets, $X_{N+1}$ is the final output results, $X_{i+1}$ is the input of $C_{i+1}$; for layer $i$, $bs$ is the batch size, $h_{i}$ and $w_{i}$ are the height and width of each feature map, $in_{i}$ is the input channel, $out_{i}$ is the output channel (filter number of $C_{i}$). We convert $X_{i}$ from the original shape to $\hat{X_{i}}$, $(bs\times h_{i}\times w_{i},in_{i})$, i.e., $(d_{i},in_{i})$. Thus, the shape of $\hat{X_{i+1}}$ is $(bs\times h_{i+1}\times w_{i+1},out_{i})$, i.e., $(d_{i+1},out_{i})$. Therefore, we use $\beta$ to denote the linkage matrix of $C_{i}$ among $X_{i}$ and $X_{i+1}$, $\beta_{j}$ denotes the $j$-th column $\beta$.
3.2 Structured Lasso
3.2.1 Information Bottleneck Function and Gram matrix
@@
For each layer $C_{i}$, we optimize on the Information Bottleneck (IB) formular Tishby et al. (2000) as:
$\min\limits_{\tilde{X_{i+1}}}-I(\tilde{X_{i+1}};X_{i})+\gamma I(X_{i+1};\tilde{X_{i+1}})$ (1)
@@
where $\tilde{X_{i+1}}$ denotes the pruned ouput features and $\gamma$ is a positive Largrange multiplier. Output feature maps are derived from input feature maps through filters and corresponding convolution operations. As such, important filters result in output feature maps that contain more information from the corresponding input feature maps. To better utilize the IB theory, we first reshaped $X_{i}$ and $X_{i+1}$ into two-dimensional matrices, denoted as $\hat{X_{i}}$ and $\hat{X_{i+1}}$. Therefore, the problem of pruning filters is transformed into the task of exploring to preserve output feature channels that contain the most statistical information with respect to the input feature maps using the Information Bottleneck approach.
$\min\limits_{\tilde{X_{i+1}}}-I(\tilde{X_{i+1}};\hat{X_{i}})+\gamma I(\hat{X_{i+1}};\tilde{X_{i+1}})$ (2)
@@
Gram matrix allows for the use of more flexible kernel functions, enabling linear operations in high-dimensional spaces, thereby reducing computational complexity and better considering the geometric structure between feature maps. To account for class-wise information expansion between feature maps, we employ Gram matrix. We transform $\hat{X_{i}}$ and $\hat{X_{i+1}}$ into centered Gram matrices $X_{i}^{g}$ and $X_{i+1}^{g}$. Therefore, we have the optimization formula as:
$\min\limits_{\tilde{X_{i+1}}}-I(\tilde{X_{i+1}};X_{i}^{g})+\gamma I(X_{i+1}^{g};\tilde{X_{i+1}})$ (3)
@@
where $X_{i}^{g}=\Psi X_{i}^{k}\Psi$, $\Psi=I_{bs}-\frac{1}{bs}1_{bs}1_{bs}^{T}$, $I_{bs}$ is a bs-dimensional identity matrix, $1_{bs}$ is a bs-dimensional vector consisting entirely of ones. The element in row $a$ and column $b$ of $X_{i}^{k}$ is equal to the inner product of $a$-th feature map and $b$-th feature map of $X_{i}$. Therefore, we have $X_{i}^{g}\in R^{bs\times bs,in_{i}},X_{i+1}^{g}\in R^{bs\times bs,out_{i}},\beta\in R^{in_{i},out_{i}}$.
@@
The first step of reshaping is intended to capture the relationship between input features and output features from the channel perspective. The second step is to utilize Gram matrix. In practice, we can directly transform $X_{i}$ to $X_{i}^{g}$.
3.2.2 Graph-structured Lasso and sGLP-IB
@@
The computation of the IB method is often considered computationally infeasible due to the significant overhead introduced by density estimation in high-dimensional spaces Ma et al. (2020). We are able to use the Lasso-based methods to get the nonlinear dependencies between the input and output of each layer Li et al. (2016); Guo et al. (2023).
$\beta=\min\limits_{\beta}\frac{1}{2}||X_{i+1}^{g}-X_{i}^{g}\beta||_{F}^{2}+\Phi(\beta)$ (4)
@@
where $||\cdot||_{F}$ denotes the Frobenius norm, and $\Phi$ is determined by structured lasso. Due to the fact that filters generally contain class-wise information Hu et al. (2016). Filters are tailored to different classes, meaning they also possess distinct statistical properties. As a result, it is possible to form corresponding subgroups based on class information. Thus, the output feature maps also conceal class-wise information due to the filters. This results in a complex relationship between output feature maps, especially in terms of statistical subgroups. Therefore, we introduce graph-structured lasso from structured lasso to better model this relationship.
@@
Given $\beta$ for $C_{i}$, we have the node sets $V$, $\beta_{1},\beta_{2},...\beta_{out_{i}}$} that forms the graph, where $\beta_{j}^{r}$ denotes element in $j$-th column and $r$-th row of $\beta$. We have weight edge set $E$ of the graph by computing pairwise Pearson correlations and linking pair of nodes when their correlation exceeds a threshold $th$, where the weight of each edge indicates correlation degree. Let $f_{lm}$ represent the weight of edge $e=(l,m)\in E$, which quantifies the correlation between element $l$ and $m$. With these, we define graph-structured lasso as:
$\Phi(\beta)=\lambda||\beta||_{1}+\mu\sum_{e=(l,m)\in E}|f_{lm}|\sum_{j=1}^{in_{i}}|\beta_{l}^{j}-sign(f_{lm})\beta_{m}^{j}|$ (5)
@@
where $\lambda$ is the hyperparameter for the sparsity and $\mu$ regulates inner class-wise dependency. Substituting (5) into (4) yields our sparse graph-structured lasso pruning with Information Bottleneck (sGLP-IB).
@@
As shown in the $\beta$ matrix in Figure 1, each column of $\beta$ represents the relatedness between all feature maps of this batch for a specific output channel and the current layer’s input feature maps. Our method can consider output feature maps and $\beta$, forming subgroups according to the classes. Within each subgroup, we identify and select the filters that contribute the most. The larger the value of elements in $\beta$, the stronger the correlation. Therefore, we can determine whether to retain a filter based on the number of non-zero values in its corresponding column.
3.2.3 Tree-Guided Lasso and sTLP-IB
@@
We also apply the state-of-the-art tree-guided lasso penalty to model the more closely related class-wise information and then form the tree structures. The tree numbers and the overlap among groups are determined by the hierarchical clustering tree due to class information. We have:
$\Phi(\beta)=\lambda\sum_{j}W_{j}(v_{root})=\lambda\sum_{j}\sum_{v\in V}w_{ve}||\beta_{j}^{G_{v}}||_{2}$ (6)

### v2

# Enhanced Structured Lasso Pruning with Class-wise Information

@@ Abstract
Modern applications require lightweight neural network models. Most existing neural network pruning methods focus on removing unimportant filters; however, these may result in the loss of statistical information after pruning due to failing to consider the class-wise information. In this paper, we employ the structured lasso from the perspective of utilizing precise class-wise information for model pruning with the help of Information Bottleneck theory, which guides us to ensure the retention of statistical information before and after pruning. With these techniques, we propose two novel adaptive network pruning schemes in parallel: sparse graph-structured lasso pruning with Information Bottleneck (sGLP-IB) and sparse tree-guided lasso pruning with Information Bottleneck (sTLP-IB). The key component is that we prune the model filters utilizing sGLP-IB and sTLP-IB with more precise structured class-wise relatedness. Compared to multiple state-of-the-art methods, our approaches achieve the best performance across three datasets and six model structures on extensive experiments. For example, with the VGG16 model based on the CIFAR-10 dataset, we can reduce the parameters by 85%, decrease the FLOPs by 61%, and maintain an accuracy of 94.10% (0.14% better than the original). For large-scale ImageNet, we can reduce the parameters by 55% while keeping the accuracy at 76.12% (only drop 0.03%) using the ResNet architecture. In summary, we succeed in reducing the model size and computational resource usage while maintaining the effectiveness of accuracy.
Introduction
@@
Deep Learning (Krizhevsky, Sutskever, and Hinton 2012), especially CNN-based methods (Simonyan and Zisserman 2014; Szegedy et al. 2015; He et al. 2016), has proven its effectiveness in tasks in numerous applications, such as multimodal fusion and vision. Recently, researchers have been attempting to deploy these neural network models on resource-constrained devices, like smartphones, for users’ real-time analysis of multi-modal information (Liu et al. 2024c; Yu et al. 2024). However, the number of these network parameters is often too large, making the deployment difficult. For instance, the VGGNet-16 model has parameters exceeding 500MB (Simonyan and Zisserman 2014), which introduces a heavy memory burden and computation overhead. Current researchers aim to explore methods that help deploy models on resource-constrained devices.
@@
Fortunately, scientists have discovered that the parameters are usually redundant and the key weights of neural networks exhibit sparsity characteristics (He, Zhang, and Sun 2017). As a result, a significant amount of work has focused on pruning unimportant channels to reduce the size of the model, which reduces the model’s FLOPs and lowers computational overhead as well. The common methods can be primarily divided into two main categories: unstructured pruning and structured pruning (Shao and Shin 2022). Unstructured pruning mainly focuses on weight pruning (Ding et al. 2019) and N:M sparsity (Zhang et al. 2022). However, both of these methods require hardware support, which makes them less scalable. Therefore, this paper primarily focuses on structured pruning, which benefits from removing unimportant filters within the CNNs and is consistent with the focus of most current pruning research.
@@
To avoid the limitations of hand-crafted approaches, most structured pruning methods utilize statistical information (Zhang, Gao, and Huang 2021; Shao and Shin 2022) to analyze filters insides CNNs for channel selection, with the aim of minimizing the loss before and after pruning and achieving better accuracy results. However, while most existing methods overlook class-wise information (Hu et al. 2016), some techniques do incorporate inter-class information (Hemmat, San Miguel, and Davoodi 2020; Hu et al. 2023; Ma et al. 2023); however, they fail to model the finer-grained class relatedness, resulting in a less precise pruning scheme and a consequent drop in accuracy. Furthermore, an ideal pruning method should also include an adaptive approach (Guo et al. 2023; Xie et al. 2024) to control the sparsity ratio according to specific requirements without the need to manually adjust hyperparameters.
@@
Based on these, we thus employ the precise Information Bottleneck (IB) (Tishby, Pereira, and Bialek 2000) theory perspective to model statistical information of feature maps of CNN for channel selection, aiming to minimize the loss from pruning. We also utilize structured lasso techniques that account for sparsity, such as graph-structured lasso (Liu et al. 2024b) and tree-guided lasso (Kim and Xing 2009; Liu et al. 2024a), which can also consider the relatedness in feature maps and form corresponding sub-groups and sub-trees. This approach helps to incorporate more class-wise information compared to previous methods. Consequently, we propose two novel methods: Sparse Graph-structured Lasso Pruning with Information Bottleneck (sGLP-IB) and Sparse Tree-guided Lasso Pruning with Information Bottleneck (sTLP-IB). Our methods perform pruning at the layer level, ensuring more precise pruning filters and preventing cumulative errors layer by layer. Based on this, our achievements have already been surprising as we are the first to integrate the precise structured class-wise information into pruning. Further, we (1) combine Gram matrix (Lanckriet et al. 2002) to reduce the dimensions and accelerate computation without compromising accuracy, (2) propose an adaptive method to ensure that our method can be automatically suitable for varying sparsity ratio requirements rather than setting hyperparameters manually based on the expert experiences, (3) use novel optimization method which has polynomial time complexity with guaranteed convergence rate. We conduct extensive experiments to evaluate the effectiveness.
Our main contributions are summarized as follows:
@@
  • We are the first to discover pruning with structured lasso from the information bottleneck perspective via precise structured class-wise information. By enhancing techniques such as graph-structured lasso and tree-guided lasso with information bottleneck, we propose two novel methods for model pruning: sGLP-IB and sTLP-IB.
@@
  • We have optimized our proposed methods, sGLP-IB and sTLP-IB, to ensure faster computation using Gram matrix and adaptively address different model sizes or FLOPs requirements with our proposed adaptive algorithm. These optimizations still guarantee convergence and polynomial computation time complexity using the proximal gradient descent method.
@@
  • We conduct comparative experiments against numerous baselines using three widely used datasets, CIFAR-10, CIFAR-100 and ImageNet. We tested multiple variants of VGGNet, ResNet, and GoogleNet. The experimental results demonstrate the effectiveness of our methods across three metrics: accuracy, memory size, and FLOPs. They both achieve state-of-the-art results, while sTLP-IB demonstrates a stronger ability to learn class information compared to sGLP-IB. Extensive ablation studies further confirm the robustness of our methods and the potential of utilizing structured class-wise information for model pruning.
Related Works
Structured Pruning
@@
MPF (He et al. 2019) considers the geometric distance between filters and neighboring filters to guide the pruning process. CPMC (Yan et al. 2021) leverages relationships between filters in the current layer and subsequent layers for pruning decisions. HRank (Lin et al. 2020) compresses models by constructing low-rank matrices using information from the same batch. CPGMI (Lee et al. 2020) leverages gradients of mutual information to measure the importance of channels to prune the less important ones. SCOP (Tang et al. 2020) reduces filters based on their deviation from the expected network behavior. GDP (Guo et al. 2021) employs gates with differentiable polarization, learning whether gates are zero during training for pruning purposes. EEMC (Zhang, Gao, and Huang 2021) uses a multivariate Bernoulli distribution along with stochastic gradient Hamiltonian Monte Carlo for pruning. SRR-GR (Wang, Li, and Wang 2021) identifies redundant structures within CNNs. FTWT (Elkerdawy et al. 2022) predicts pruning strategies using a self-supervised mask. DECORE (Alwani, Madhavan, and Wang 2022) assigns an agent to each filter and uses lightweight learning to decide whether to keep or discard each filter. EPruner (Lin et al. 2021) employs Affinity Propagation for efficient pruning. DPFPS (Ruan et al. 2021) directly learns the network with structural sparsity for pruning. CC (Li et al. 2021) combines tensor decomposition for filter pruning. NPPM (Gao et al. 2021) trains the network to predict the performance of the pruned model, guiding the pruning process. LRF-60 (Joo et al. 2021) uses Linearly Replaceable Filters to aid pruning decisions. AutoBot (Castells and Yeom 2021) tests each filter one-by-one to ensure pruning precision. PGMPF (Cai et al. 2022) utilizes prior masks as the basis for pruning masks. DLRFC (He et al. 2022) uses Receptive Field Criterion to measure filter importance. FSM (Duan et al. 2022) aggregates feature and filter information to evaluate their relevance for shifting. Random (Li et al. 2022) employs random search strategies for pruning. The study (Hussien et al. 2024) employs activation statistics, foundations in information theory, and statistical analysis with designed regularizations for pruning. CSD (Xie et al. 2024) uses channel spatial dependability metric and leverages feature characteristics for pruning. However, the statistical learning information considered by these methods may not be precise enough.
Information Bottleneck and Lasso Principle
@@
The Information Bottleneck (IB) (Tishby, Pereira, and Bialek 2000) method extracts the most relevant output information by considering the input. Aside from the previous mentioned methods, the works (Tishby and Zaslavsky 2015; Dai, Zhu, and Wipf 2018) were the first to use IB for compressing CNNs. FPGM (He et al. 2018) employs the geometric median for pruning. Li et al. (Li et al. 2016) illustrates the correspondence between Lasso and IB for pruning. The approach in (He, Zhang, and Sun 2017) utilizes a two-step optimization process incorporating Lasso for pruning. Papers (Scardapane et al. 2016; Rajaraman et al. 2023) leverage the $l_{2}$ lasso (e.g., group lasso) for pruning, with analyzed complexity (Rajaraman et al. 2023). APIB (Guo et al. 2023) integrates IB with the Hilbert-Schmidt Independence Criterion Lasso for exploration. The paper (Sakamoto and Sato 2024) also shows through analysis that layer-level pruning, based on the Hilbert-Schmidt independence criterion, is preferable to end-to-end pruning.
@@
Motivation: However, none of these methods account for class-wise information. To address this, from the information bottleneck perspective, we employ state-of-the-art techniques like graph-structured lasso (Liu et al. 2024b) and tree-guided lasso (Liu et al. 2024a) to aggregate structured class-wise information effectively to propose our novel methods and prune filters inside CNNs. The following section explains how we combine Information Bottleneck with structured Lasso as well as reducing computational complexity without compromising accuracy, and how we integrate our adaptive method with our novel methods to control the sparsity.
Figure 1: The overview of our methods. To utilize information bottleneck theory, ① reshape the input and output feature maps. ② transforms the reshaped feature maps with Gram matrix. Then, we utilize the structured lasso to account for the structured class-wise information to get the linkage matrix and prune the filters.
Methodology
Design Overview
@@
The overview of our methods is shown in Fig. 1. To avoid the imprecision that end-to-end methods might introduce, we perform layer-wise pruning independently. We reshape each layer’s input and output feature map tensors into two-dimensional matrices, enabling us to apply the information bottleneck theory to determine which filters can be pruned. To further reduce computational complexity, capture internal geometric structures, and utilize more flexible kernel functions, we employ Gram matrix (Lanckriet et al. 2002). Since lasso aligns well with the information bottleneck theory (Tishby and Zaslavsky 2015; Dai, Zhu, and Wipf 2018), we use structured lasso for solving this problem. This approach also takes into account structured class-wise information in the feature maps (as well as filters), resulting in a sparse linkage map between layers’ feature maps for pruning.
@@
Math Notations: In this paper, given a CNN model $M$ with $N$ layers, we let $\bm{C}_{i}$ denote the $i$-th layer of $M$. The input feature map of $\bm{C}_{i}$ is $\bm{X}_{i}$ with the shape $(bs,h_{i},w_{i},in_{i})$; the output feature map of $\bm{C}_{i}$ is $\bm{X}_{i+1}$ with the shape $(bs,h_{i+1},w_{i+1},out_{i})$, where $i\in\{1,2,...,N\}$, $\bm{X}_{1}$ is the input sample batch from the datasets, $\bm{X}_{N+1}$ is the final output results, $\bm{X}_{i+1}$ is the input of $\bm{C}_{i+1}$; for layer $i$, bsbs is the batch size, $h_{i}$ and $w_{i}$ are the height and width of each feature map, $in_{i}$ is the input channel, $out_{i}$ is the output channel (filter number of $\bm{C}_{i}$). We convert $\bm{X}_{i}$ from the original shape to $\hat{\bm{X}_{i}}$, $(bs\times h_{i}\times w_{i},in_{i})$, i.e., $(d_{i},in_{i})$. Thus, the shape of $\hat{\bm{X}_{i+1}}$ is $(bs\times h_{i+1}\times w_{i+1},out_{i})$, i.e., $(d_{i+1},out_{i})$. Therefore, we use $\bm{\beta}$ to denote the linkage matrix of $\bm{C}_{i}$ among $\bm{X}_{i}$ and $\bm{X}_{i+1}$, $\bm{\beta}_{j}$ denotes the $j$-th column $\bm{\beta}$.
Structured Lasso
Information Bottleneck Function and Gram matrix
@@
For each layer $\bm{C}_{i}$, we optimize on the Information Bottleneck (IB) formular (Tishby, Pereira, and Bialek 2000) as:
[…]
@@
where $\tilde{\bm{X}_{i+1}}$ denotes the pruned ouput features and $\gamma$ is a positive Largrange multiplier. Output feature maps are derived from input feature maps through filters and corresponding convolution operations. As such, important filters result in output feature maps that contain more information from the corresponding input feature maps. To better utilize the IB theory, we first reshaped $\bm{X}_{i}$ and $\bm{X}_{i+1}$ into two-dimensional matrices, denoted as $\hat{\bm{X}_{i}}$ and $\hat{\bm{X}_{i+1}}$. Therefore, the problem of pruning filters is transformed into the task of exploring to preserve output feature channels that contain the most statistical information with respect to the input feature maps using the Information Bottleneck approach.
[…]
@@
Note that Gram matrix allows for the use of more flexible kernel functions without the loss of information, enabling linear operations in high-dimensional spaces, thereby reducing computational complexity and better considering the geometric structure between feature maps. Thus, to better account for structured class-wise information between feature maps and reduce computation complexity, we employ Gram matrix. We transform $\hat{\bm{X}_{i}}$ and $\hat{\bm{X}_{i+1}}$ into centered Gram matrices $\bm{X}_{i}^{g}$ and $\bm{X}_{i+1}^{g}$. Therefore, we have the optimization formula:
[…]
@@
where $\bm{X}_{i}^{g}=\bm{\Psi}\bm{X}_{i}^{k}\bm{\Psi}$, $\bm{\Psi}=I_{bs}-\frac{1}{bs}1_{bs}1_{bs}^{T}$, $I_{bs}$ is a bs-dimensional identity matrix, $1_{bs}$ is a bs-dimensional vector consisting entirely of ones. The element in row $a$ and column $b$ of $\bm{X}_{i}^{k}$ is equal to the inner product of $a$-th feature map and $b$-th feature map of $\bm{X}_{i}$. Therefore, we have $\bm{X}_{i}^{g}\in R^{bs\times bs,in_{i}},\bm{X}_{i+1}^{g}\in\bm{R}^{bs\times bs,out_{i}},\bm{\beta}\in\bm{R}^{in_{i},out_{i}}$.
@@
The first step of reshaping is intended to capture the relationship between input features and output features from the channel perspective. The second step is to utilize Gram matrix. In practice, we can directly transform $\bm{X}_{i}$ to $\bm{X}_{i}^{g}$.
Graph-structured Lasso and sGLP-IB
@@
The computation of the IB method is often considered computationally infeasible due to the significant overhead introduced by density estimation in high-dimensional spaces (Ma, Lewis, and Kleijn 2020). We are able to use the Lasso-based methods to get the nonlinear dependencies between the input and output of each layer (Li et al. 2016; Guo et al. 2023).
[…]
@@
where $||\cdot||_{F}$ denotes the Frobenius norm, and $\Phi$ is determined by structured lasso. Due to the fact that filters generally contain class-wise information (Hu et al. 2016). Filters are tailored to different classes, meaning they also possess distinct statistical properties. As a result, it is possible to form corresponding subgroups based on class information. Thus, the output feature maps also conceal structured class-wise information due to the filters. This results in a complex relationship between output feature maps, especially in terms of statistical subgroups. Therefore, we introduce graph-structured lasso from structured lasso to better model this relationship.
@@
Given $\bm{\beta}$ for $\bm{C}_{i}$, we have the node sets $V$, $\bm{\beta}_{1},\bm{\beta}_{2},...\bm{\beta}_{out_{i}}$} that forms the graph, where $\bm{\beta}_{j}^{r}$ denotes element in $j$-th column and $r$-th row of $\bm{\beta}$. We have weight edge set $E$ of the graph by computing pairwise Pearson correlations and linking pair of nodes when their correlation exceeds a threshold $th$, where the weight of each edge indicates correlation degree. Let $f_{lm}$ represent the weight of edge $e=(l,m)\in E$, which quantifies the correlation between element $l$ and $m$. With these, we define graph-structured lasso as:
[…]
@@
where $\lambda$ is the hyperparameter for the sparsity and $\mu$ regulates inner class-wise dependency. Substituting (5) into (4) yields our sparse graph-structured lasso pruning with Information Bottleneck (sGLP-IB) method.
@@
As shown in the $\bm{\beta}$ matrix in Fig. 1, each column of $\bm{\beta}$ represents the relatedness between all feature maps of this batch for a specific output channel and the current layer’s input feature maps. Our method can consider output feature maps and $\bm{\beta}$, forming subgroups according to the classes. Within each subgroup, we identify and select the filters that contribute the most. The larger the value of elements in $\bm{\beta}$, the stronger the correlation. Therefore, we can determine whether to retain a filter based on the number of non-zero values in its corresponding column.
Tree-Guided Lasso and sTLP-IB
@@
We also apply the state-of-the-art tree-guided lasso penalty to model the more closely related structured class-wise information and then form the tree structures. The tree numbers and the overlap among groups are determined by the hierarchical clustering tree due to class information. We have:
[…]